OpenStudy (anonymous):

which point lies in the solution set for the following system of inequalities? x-2y>3 x+2y<6

OpenStudy (anonymous):

first ignore the initial inequalities, i.e: x-2y=3; x+2y=6 solve these: x=3+2y => 3+2y+2y=6 => 4y=3 => y=3/4 => x=4.5 now you have the critical points, test both inequalities to see which side of the critical point holds

OpenStudy (anonymous):

(3,4) @epicfail

OpenStudy (anonymous):

(3,4) is not the solution, from 1st equation: (3-2(4)) does not equal 6

OpenStudy (anonymous):

just do the four combinations: x>4.5, y>.75 x<4.5, y>.75 x>4.5, y<.75 x<4.5, y<.75 and the one which hold is correct

OpenStudy (anonymous):

(4,-2)

OpenStudy (anonymous):

@epicfail